OpenStudy (anonymous):

find the product 5 1/4 * 1 1/2

OpenStudy (lgbasallote):

change the first number to improper fraction then proceed to how you multiply fractions

OpenStudy (anonymous):

9/4 * 3/2 = 27/8

OpenStudy (callisto):

@pratu043 i think 5 1/4 is (5x4+1)/4 = 21/4 ...

OpenStudy (anonymous):

Oh no stupid mistake ... my exams got over today and I'm in no mood to work out sums ...

OpenStudy (anonymous):

Thanks for pointing out my mistake Callisto.

OpenStudy (anonymous):

Which one 6 1/8 5 /1/4 5 1/8 7 7/8

OpenStudy (arnavguddu):

5 1/4 * 1 1/2 = 21/4 * 3/2 = 63/8 = 7 7/8

OpenStudy (anonymous):

It is 7 /78